Sol Summary – December 20th

Crew 185 Sol Summary Report  20 Dec 2017
Sol 2
Crew 185’s First EVA
David Murray
Mission Status:
Sol Activity Summary: The crew woke up, ate breakfast together, and attached our patches to our flight suits. Then, with the help of the entire crew, Tibo donned the X-1 space suit and headed out with Arno and I, in EVA suits, to complete the daily engineering check. Immediately following this, Tibo changed into the EVA suit to join Ilaria and I for our first EVA. We walked for about 3 km and explored the surrounding terrain. When we returned, we continued our custom of eating meals as a team and sat down for lunch. Following our meal, I gave the crew a quick tour and summary of my research and work in the Greenhab. We then returned to the Hab to take part in John’s anesthesia training where we learned how to identify and anesthetize the sciatic nerve on a gel block with the help of ultrasound imaging.
Look Ahead Plan: We plan to go on our second EVA tomorrow. It should be the same as the first EVA but with different crew members.
Anomalies in work:  None
Weather: Sunny, warm
Crew Physical Status: Healthy
EVA: 3 km EVA with no vehicles to explore the surrounding terrain and get used to the EVA suits.
Reports to be file: None
Support Requested: None

Sol Summary – December 19th

Crew 185 Daily Summary Report  19/12/2017
MDRS Daily Summary Report for Sol 1
First Day of Sim
Mission Status:  Things are running smoothly.
Sol Activity Summary: We started our simulation this morning. We began our day by eating breakfast as a team and followed our meal with a mindfulness meditation session. After that, three of us put on our simulation space suits for the first time and left the hab to do an engineering check. Two crew members stayed behind for communication. When the crew returned to the hab, we ate lunch, assigned daily tasks for each crew member, and had time to work on individual projects.
Look Ahead Plan: Tomorrow, we will have our first EVA and test the X-1 space suit for the first time.
Anomalies in work:  The simulation space suit helmet fogs up easily and can drastically limit your field of vision.
David Murray crew greenhab officer and biologist
Weather: Sunny, warm
Crew Physical Status: Healthy
EVA: No EVA today
Reports to be file: EVA request
Support Requested: None

Sol Summary – December 18th

Today has been a busy day for the Crew 185. As it was our last day before entering simulation, we needed to prepare the hab in the best possible way and to fix as many things as possible before closing our doors.

We decided to work in two groups. Ilaria and David took the car to buy a new battery for the generator, soil for the greenhab and food. Arno, Thibault and John set the laboratory with the 3D printer, the medical supplies, the X-1 spacesuit prototype, etc…

This afternoon, we repaired the generator by replacing the battery and changing a part of the cooling liquid and the oil. It is working better at the moment when I write these lines! Hopefully, it will keep being in a good shape.

We decided that we would begin the simulation tomorrow morning when waking up. Mars, here we come!

The Crew 185

Sol Summary – December 15th

MDRS Sol Summary Report for Sol 14

Summary Title: Signing off!

Mission Status:

It’s been a please working with all of you, and thanks for your help.

It was a fitting send off to our team that our day started with a
generator failure. Upon our morning rounds after waking up it was
noticed the Generator had failed with an ‘Underspeed’ Fault. We were
happy to see that SOC was holding steady at 100% (thank you solar
arrays!) but this was a worrying sign, especially as we didn’t want to
hand over a crippled ship to the oncoming crew.

Investigation into the generator showed that the loose connections
that were repaired previously were still holding strong, so there was
no immediate cause of the generator fault. After attempting to
restart the generator a low battery voltage was observed (reading
7.6volts, supposed to be 9.6 volts). It appears that after the
generator failed it drained the battery down displaying the fault
message on its status plate. After a jumpstart from our friendly
neighbors we were able to get the generator back up and running. This
issue is likely to re-occur as we are still not sure of the original
cause. Our next step will be to Remove and Replace the generator
batteries, fingers crossed this fixes the problem. Best of luck to
future crews in continuing to troubleshoot this issue.

Besides that morning excitement the day was one of cleaning and
packing, with a short break in the afternoon to do a short EVA around
the HAB for some last minute filming to get the shots we’ve missed so
far this mission. We hope to leave this place better than we found
it.

This report comes to you with the assistance of Mission Support as our
Internet allowance has run out. Sometime today the same laptop that
killed our bandwidth before had WiFi turned On, after which it sucked
our bandwidth dry again. Still don’t know why this particular laptop
uses bandwidth at such a prodigious rate, but we’ve turned it off
again and comms should be returned to normal.

Hope to work with you all again in the future.

Ad Astra Per Ardua,

Crew 184.

Sol Activity Summary:

1. Filming EVA

2. Cleaning / Packing

4. Reports and general HAB maintenance / checks

6. Generator Failure and Jumpstart

Look Ahead Plan:

1. Crew Handover Training

Anomalies in work:

1. HAB Leak from under EVA door continuing (per Mission Control
leak is expected and has been present for months, presumed to be from
Kitchen piping)

2. Generator Underspeed Fault

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.

EVA: None today.

Reports to be filed:

Greenhab

Operations

Journalist

Astronomy

Sol Summary

Mission Summary

Support Requested:

1. None

Sol Summary – December 14th

Summary Title:  Wrapping Up!

Mission Status:

Today marks the end of our scientific mission and the completion of our three remaining scientific objectives.

Our martian day sleep study ends tonight with a 4am bedtime, 12pm wake up.  Our team prepped extensively for this and arranged numerous protocols to handle crew fatigue and performance drop off due to tiredness.  Luckily we haven’t had to use any!  Our team has remained in good spirits and been able to keep to our sleep schedule well throughout our mission, with the occasional controlled nap being used to support a flagging crewmember!  We’ve focused on mentally demanding activities and music to keep the crew focused and alert in the evenings to allow us to hit our sleep targets.

Today we successfully completed our last Matryoshka objective!  This was the farthest we’ve had to travel for an EVA, which is why we left it till last to capitalize on the experience we’ve gained in the past two weeks.  We took the Deimos Rover and an ATV to ensure we had sufficient range to meet our objectives and departed from the HAB at around 1pm, arriving at our objective at around 2pm in the Blue Hills area.  Our goal was to get samples from the ridgeline as it washed out into the plane of dirt to try to identify different materials in the mixing area between the cliffs and plains.  We successfully took samples from this area and several isolated hilltops in the area and packed up from or last EVA to return home.

We weren’t going to get off easy however, as that hasn’t been our type of mission.  Due to user error when we started home the parking break was left on for ~ 5 minutes, which resulted in a precipitous power drop off from ~ 80% to 40%.  After the problem was noticed and corrected we discussed options.  The Blue Hills area is actually a fairly close walking distance from the HAB, but a long driving distance, so the decision was weather to leave the rover and walk back to the HAB or to take the long way around on the roads and trust the Rover to make it.  After assessing the fact that the Rover utilized ~ 20% battery to get to this location, had 40% remaining, and that we had a functioning ATV to get help if we broke down on the way we decided we had high confidence in getting the Rover back to the HAB and started the long drive back.  It made for an anxious drive back and frequent map checks to ensure we were going the right way but we managed to successfully coax Deimos back to the HAB just as the low battery light turned on at 20% SOC.  Another successful, and exciting, EVA!

Today we completed our Anesthesia runs ‘for score’.  Under the supervision and evaluation of our resident doctor two crewmembers at a time formed a team and had to evaluate a test mold he has created to mimic a human leg.  Inside the mold the team must use the instructional videos and procedures provided to identify various components inside the mold (veins / arteries / nerves / etc. . .), identify the proper spot to apply anesthesia, then go through the process of delivering anesthesia to the leg.  Hopefully this research and experimentation makes good progress towards developing medical trainers that can be used with astronauts on a long duration mission that may not have anyone on board with medical expertise.

Unfortunately cloud cover prevented any viewing of the Geminid meteor shower last night (of course the only cloudy night was the one with the meteor shower!)  It is currently clear and we intend to try again tonight to see if we can catch the tail end of the meteor shower.

Tomorrow we focus on MDRS cleaning, packing, and transition to the next crew.  We will also be doing an EVA strictly in the environs of MDRS to get some last filming done as requested by our journalist.

Sol Activity Summary:

  1. Matryoshka EVA Site #7a
  2. Exercise
  3. General HAB Cleaning / Maintenance / Reports
  4. Anesthesia Experiment
  5. Beginning of cleanup / packing
  6. (Hopefully)  Telescope / Geminid Observation tonight due to clouds last night!

Look Ahead Plan:

  1. EVA Filming around the HAB
  2. Solar Observation / Observatory Shutdown
  3. MDRS Cleaning
  4. Lessons Learned Compilation
  5. End of Mission Food Inventory
  6. Mission Summary
  7. Exercise
  8. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.    HAB Water Leak continues from under EVA door.  It appears to be getting worse.

2.  Power System Misconfiguration ongoing.

Weather: Warm (above freezing)

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

Sol Summary

Journalist

Support Requested:  None

Sol Summary – December 13th

Summary Title:  In the final stretch (and clouds)

Mission Status:

We are getting close to the end of our mission, and the focus of the crew is maximizing the time we have left to complete our science objectives.

Today in the Matryoshka EVA series. Crewmembers Trisha Randazzo, Joshua Hunt, and Willie Schumann traveled to “Matrysoshka Site 6” which was located in the far north at the begging of Lith Canyon. After a long drive to the site with

We attempted to do solar observation tonight but were unfortunately stymied by extensive cloud cover all day long today.

This evening our team will focus on Anesthesia Experiment runs.  Once complete our team is currently hoping the clouds clear to allow us to watch the Geminid meteor shower which peaks tonight.  Shannon also allowed us to assemble her old telescope for use tonight.  Fingers crossed the clouds clear!

We preparing for our last ‘science day’ on Thursday, as we finish up both Matryoshka, Bricks, and Anesthesia experiments.  Friday is being held in reserve and is also allocated to cleaning the HAB and writing up lessons learned for the next crew.  The end is fast approaching!

Thanks for your help!

Sol Activity Summary:

  1. Matryoshka EVA Site #6
  2. ‘Bricks’ communication experiment
  3. Exercise
  4. General HAB Cleaning / Maintenance / Reports
  5. Telescope Assembly
  6. Anesthesia Experiment Runs this evening
  7. (Hopefully)  Geminid Observation!

Look Ahead Plan:

  1. Matryoshka EVA Site #7a
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Anesthesia Experiment Runs
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.    HAB Leak from under EVA door.  No crew has showered yet today and a significant leak was observed by returning EVA crew.  It is likely this leak is coming from the Kitchen as the dishes were washed immediately prior to the crew returning.

2.  Power System Misconfiguration ongoing.

Weather: Mild and Cloudy

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

Support Requested:

1.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
    16. 1 Can Strawberries
    17. 1 Box Granola
    18. 1 Bottle Soy Sauce

Sol Summary – December 12th

Summary Title:  Getting back on routine (and our first EVA failure!).

Mission Status:

Today was a successful day and will hopefully  mark out the routine we will follow until the end of our mission.  The generator has been completely functional now for two days since our repairs over the weekend which is thankful news, our power problems seem to have abated themselves for the mission (though I may be jinxing myself. . . .).  In addition, the weather has been getting warmer which is a welcome relief and allows me to finally stop wearing my thermals, which is a welcome relief as they could sorely use a wash!

We resumed this week’s round of Matryoshka EVAs with a successful visit and sample collection to Site #5.  Just two more sites to go and we will have completed our most important EVA objective of our mission.  We managed to collect several soil and rock samples from what appeared to look like river channels. We saw shiny glinting on the hills too and found some shale like specimens for analysis. All in all a successful EVA! One thing that has come about is the hazard of the strong sunlight in your face reducing visibility, as there is no sun visor on the helmet and sunglasses are not recommended. Luckily it caused no issues on this EVA, but could potentially be dangerous if sunglasses are to remain forbidden and visors are not added to EVA helmets.

During the EVA return our EVA crew encountered a failure with the Deimos Rover.

Deimos left the HAB at SOC 100% and was driven for ~ 45 minutes to the site with no issues.

When returning from EVA site Deimos started experiencing significant mechanical drag that both lowered the achievable rover speed and increased battery drain.  This problem persisted for roughly 30 minutes until the Rover was able to return to base.  From the onset of the problem it was roughly steady until the turn was made off of cowdung road to return to the HAB.  At that time Rover performance precipitously dropped and was able to achieve little more than walking speed.  Steering was also noticed to be stiff during the return from EVA site.

The EVA crew reports that from a driving perspective it felt like the parking break was on, though visual inspection showed this was not actually the case.  The root cause of the issue was not determined though we will do an inspection during tomorrows EVA.  We suspect the issue may be due to debris in Deimos’s drive train.

At the end of the EVA Deimos was at 29% which is unusual given Deimos’s long range and is likely related to the experienced ‘drag’ when driving Deimos.

For the folks left behind today we were primarily occupied with Solar observation which showed numerous impressive features that made observing a treat.  The sun is waking up!  Hopefully it stays active for the next few days and into the next mission.

We are entering the period of our mission where our Martian Sleep Schedule has moved our days ‘out of phase’ with Earth normal.  Tonight we will be going to sleep at 2:40am and waking up at 10:40am.  We have been primarily occupying our nights with cleaning, Bricks, Anesthesia training, and general maintenance to prepare ourselves to fully utilize the next day’s daylight hours.  Our team has adjusted to the schedule well so far with crew taking naps as needed to stay on schedule.

Thanks for your help!

Sol Activity Summary:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Exercise
  5. General HAB Cleaning / Maintenance / Reports
  6. EVA Suit Inspection / Maintenance

Look Ahead Plan:

  1. Matryoshka EVA Site #6
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Anesthesia Experiment Runs
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports
  7. Deimos Troubleshooting


Anomalies in work:

1.    HAB Leak from under EVA door (not noticeable today)

2.  Power System Misconfiguration ongoing.

3.  Deimos Rover (see above)

Weather: Mild

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

HSO

Support Requested:

1.    ATV Fuel (we are empty)

2.  Burnable / Non Burnable Trash Pickup from Rear Airlock

3.  Permission to replant seedlings into individual planters so herbs have clean soil.  See Greenhab Report for details.

4.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

5.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
    16. 1 Can Strawberries
    17. 1 Box Granola
    18. 1 Bottle Soy Sauce

Sol Summary – December 11th

Summary Title:  A nominal day!

Mission Status:

Today has been a thankfully nominal day!  Due to our error with comm’s yesterday and not getting our EVA request in on time we stayed at MDRS and concentrated on catching up on HAB maintenance and other activities that have fallen by the wayside in the last few days.

Today we completed our crew training on our ultrasound machine, which enables us to complete our practice runs with the Astronaut trainer developed by Dr. Sczepaniak in the next few days.

We were informed of some likely solar activity by our friendly neighborhood astronomer which led us to perform some highly anticipated solar observation this afternoon.  While we could not detect any prominence we did detect sunspots for the first time.  We are hoping the sun continues in this period of high activity and future solar observation yields some rewarding images.

We continued with our communication experiment and have noticed significant improvement among the crew in terms of numbers of errors and completion times in our building tasks.  We have assembled a lesson’s learned list of communication tips to allow quick and efficient communication among various types of time delay.  We will continue this experiment in various permutations until our mission ends, and then expect to hand over our experiment materials to make this a permanent installation at MDRS that crews can participate in as desired.

We also completed our CPR techniques experiment runs on all crew, analyzing the various ways that more force than body weight can be utilized during a CPR medical emergency.  It was interesting that most crew actually utilized different techniques during this experiment run, with us comparing notes to identify the optimal method.

Our crew has returned to our exercise routine today with all exercise equipment completely assembled.  This has proven to be a crew favorite and we also intend to donate this equipment to MDRS.  We have assembled a stationary exercise bike and a series of medicine balls for weights exercises.

The remainder of the day was filled with HAB maintenance, static water tank transfer, cleaning, some much needed love for our friendly plants in the Greenhab, and assembly of equipment delivered by Mission Control.  We are excited to finish out our mission and eager to get back out into the Martian environment with our 5th Matryoshka experiment site tomorrow.

Thanks for your assistance!

Sol Activity Summary:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports
  7. EVA Suit Inspection / Maintenance
  8. Washvac Assembly
  9. Power system monitoring
  10. Greenhab Planting
  11. Anesthesia Ultrasound Training
  12. Static Water Tank Transfer

Look Ahead Plan:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration ongoing.  Generator turned off ~ 10am.  SOC stayed close to 100% all day long.

3.  EVA Suit Malfunctions – See Operations Report

Weather: Cold, clear

Crew Physical Status: All crew in good physical health.

EVA: N/A

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

Support Requested:

1.    ATV Fuel

2.  Burnable / Non Burnable Trash Pickup from Rear Airlock

3.  Permission to replant seedlings into individual planters so herbs have clean soil.  See Greenhab Report for details.

4.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

3.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken

Sol Summary – December 10th

MDRS Daily Summary Report for Sol 9

Summary Title:  Return of the Sim!

Mission Status:

Yesterday and today has been been primarily ‘out of sim’ due to the numerous HAB failures, including our evacuation last night due to a loss of power.  We spent the morning today working to repair the generator and appear to have gotten it in working order, though further work is necessary to ensure it’s long term successful operation (see Shannon’s email for details).  The solar array system is still misconfigured so power problems may yet pop up during the rest of our sim but we will take them as we go and troubleshoot, as a real martian crew would.

This afternoon we took the rovers out to drain the batteries per Mission Control request, once returned to the HAB we re-entered SIM at roughly 5pm.

It’s been a rough first week, but we are looking forward to getting back into sim and finishing out our mission successfully.

Sol Activity Summary:

1.      Out of SIM maintenance work

2.     General HAB cleaning / reports

Look Ahead Plan:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration ongoing

3.  2 EVA suit fans were non-functional during HAB check today, which may be related to power issues and resulting in them not being charged.  We have connected the suits to power and will assess tomorrow.  

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.

EVA: N/A

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Support Requested:

1.      ATV Fuel

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken

Sol Summary – December 7th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Back on schedule!

Mission Status:  The past two days have been hectic due to getting in to the swing of things on Mars and the numerous failures we have encountered.  As a consequence, our schedule fell by the wayside and we were largely winging it.  Yesterday we sat down, came to a stable configuration, discussed our lessons learned from our stay so far, and prepared to jump back onto our schedule.

We are still dealing with our power system failure, but today has been a day of focus where our team is efficiently moving from task to task and completing our assigned activities.

We’ve learned today that the low battery SOC we’ve been seeing since our mission started is likely due to a misconfigured battery panel.  Mission Control is seeking help to reprogram it and return our power system to nominal.  In the meantime we will continue to minimize power usage.  Due to SOC ~ 35% it is possible we lose power tonight, our team has prepared for this eventuality and everyone is prepared to handle the cold if necessary, and all team members have light sources on hand overnight.  We have also filled reserve containers with water due the expectation that if power is lost pipes will freeze overnight.

  1. Power system is still smart enough to accept power from generator once SOC starts dropping overnight.  System Nominal!
  2. Power system swaps to generator power.  Generator Failure overnight due to the cold.  Expected loss of power overnight
  3. Power System misconfiguration prevents accepting generator power.  Expected loss of power overnight.

In the meantime, activities!

This morning we have completed Anesthesia training for all crewmembers.  On a long duration mission it is critical crews know how to treat any injuries or other problems that arise, as part of this devising methods of training for crew on a long duration mission is critical.  Our Dr. has devised a mold that simulates arteries / nerves / and veins in a leg that allows crew to practice using ultrasound equipment to identify and isolate specific body parts.  With the training complete crew will be able to complete trial runs where we see how effective this is a training device and how crew performance increases over time.

We also completed our third Matryoshka EVA to SIte #3.  We continue to collect geological samples from our target sites, though this is the first day that our crew is starting to feel the fatigue from EVA, which is hard work!.  Today’s site was exceptionally rocky which gave a very ‘martian’ feel.

Our crew is looking forward to the weekend which we have schedule as ‘light duty’ days to allow ourselves to recuperate from a very busy week!

Sol Activity Summary:

1.      Matryoshka EVA Site #3

2.      Observatory Operation

3.      Anesthesia Training

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel and submission to

6.      Exercise

7.  Brick Experiment

8.  Power System Monitoring

9.  Receipt and stowage of supply drop

10.  CPR techniques (ME-3)

11.  Media Visit Cancelled

Look Ahead Plan:

  1. Matryoshka EVA Site #4
  2. ANAM Winscat Baseline
  3. Musk Observation
  4. CPR Tecniques
  5. Brick Experiment
  6. Filming and Interviews
  7. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration

Weather: Still cold (below freezing).

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Astronomy

Sol Summary

Support Requested:

1.      Trash pickup from rear airlock (Non Burnable)

2.      Replacement ATV Fuel